# Large Language Models for Cyber Security: A Systematic Literature Review

### v2

@@ 5.1. Fine-tuning LLMs for Security Tasks
A total of 32 papers in this study applied fine-tuning techniques to address security tasks. Among them, the most popular approach is full fine-tuning, with 23 papers, accounting for 71.88% of the total. This may be because the pre-training tasks of LLMs are far from the content of the security tasks being applied, thus requiring updating all parameters of LLMs to achieve more competitive performance. partial fine-tuning is also highly regarded, with 28.12% of the papers choosing this approach to fine-tune LLMs. As shown in Table 6, full fine-tuning has a wide range of applications, including information and content security, network security, and software and system security. The most widespread domain among them is software and system security, totaling 15 papers, accounting for 65.22% of the total. A similar distribution is also observed in partial fine-tuning, with the most widely applied domain still being software and system security, totaling 7 papers, accounting for 77.78%. The applicability of fine-tuning techniques in these security tasks indicates that pre-training LLMs may not adequately address these security tasks, and updating model parameters on specific datasets is necessary to enhance effectiveness. The choice between full fine-tuning and partial fine-tuning depends on the balance between performance and efficiency considerations.
@@ 5.3. External Augmentation
External augmentation techniques have significantly boosted the effectiveness of LLMs across various security tasks, yielding competitive performance. We observed that only 26 out of the total 127 papers in LLM4Security, accounting for 20.47%, applied specific external augmentation techniques. From these studies, it is evident that external augmentation techniques have the potential to address issues such as hallucinations and high false positive rates caused by deficiencies in LLMs’ domain knowledge and task alignment. We believe that the integration of LLMs with external techniques will be a trend in the development of automated security task solutions.
RQ3 - Summary (1) We summarize the domain-specific techniques used in previous research to apply LLMs to security tasks, including prompt engineering, fine-tuning, and external augmentation. (2) Prompt engineering is the most widely used domain technique, with almost all 127 papers employing this approach. Fine-tuning techniques were used in 25.2% of the papers, while task-specific external augmentation techniques were adopted in 20.47% of the papers. (3) We categorize and discuss the fine-tuning and external augmentation techniques mentioned in these papers, and analyze their relevance to specific security tasks.
@@ 8.1. Challenges
The generalization capability of LLMs pertains to their ability to consistently and accurately execute tasks across diverse tasks, datasets, or domains beyond their training environment. Despite undergoing extensive pre-training on large datasets to acquire broad knowledge, the absence of specialized expertise can present challenges when LLMs encounter tasks beyond their pre-training scope, especially in the cybersecurity domain. As discussed in Section  4, we explored the utilization of LLMs in 21 security tasks spanning five security domains. We observed substantial variations in the context and semantics of code or documents across different domains and task specifications. To ensure LLMs demonstrate robust generalization, meticulous fine-tuning, validation, and continuous feedback loops on datasets from various security tasks are imperative. Without these measures, there’s a risk of models overfitting to their training data, thus limiting their efficacy in diverse real-world scenarios.
@@ 9. Conclusion
LLMs are making waves in the cybersecurity field, with their ability to tackle complex tasks potentially reshaping many cybersecurity practices and tools. In this comprehensive literature review, we delved into the emerging uses of LLMs in cybersecurity. Our analysis covered the different LLMs employed in security tasks, discussing their unique traits and applications (RQ1). We then explored the diverse array of security tasks where LLMs have been deployed, highlighting their practical impacts (RQ2). Additionally, we examined domain-specific techniques for applying LLMs to security tasks (RQ3). Lastly, we scrutinized the data collection and preprocessing procedures, underlining the importance of well-curated datasets in effectively applying LLMs to address security challenges (RQ4). We outlined key challenges facing LLM4Security and provided a roadmap for future research, outlining promising avenues for exploration.

### v3

@@ 5.1. Fine-tuning LLMs for Security Tasks
A total of 32 papers in this study applied fine-tuning techniques to address security tasks. Among them, the most popular approach is full fine-tuning, with 23 papers, accounting for 71.88% of the total. This may be because the pre-training tasks of LLMs are far from the content of the security tasks being applied, thus requiring updating all parameters of LLMs to achieve more competitive performance. partial fine-tuning is also highly regarded, with 28.12% of the papers choosing this approach to fine-tune LLMs. As shown in Table 7, full fine-tuning has a wide range of applications, including information and content security, network security, and software and system security. The most widespread domain among them is software and system security, totaling 16 papers, accounting for 65.57% of the total. A similar distribution is also observed in partial fine-tuning, with the most widely applied domain still being software and system security, totaling 7 papers, accounting for 77.78%. The applicability of fine-tuning techniques in these security tasks indicates that pre-training LLMs may not adequately address these security tasks, and updating model parameters on specific datasets is necessary to enhance effectiveness. The choice between full fine-tuning and partial fine-tuning depends on the balance between performance and efficiency considerations.
@@ 5.3. External Augmentation
External augmentation techniques have significantly boosted the effectiveness of LLMs across various security tasks, yielding competitive performance. We observed that only 28 out of the total 127 papers in LLM4Security, accounting for 22.05%, applied specific external augmentation techniques. From these studies, it is evident that external augmentation techniques have the potential to address issues such as hallucinations and high false positive rates caused by deficiencies in LLMs’ domain knowledge and task alignment. We believe that the integration of LLMs with external techniques will be a trend in the development of automated security task solutions.
RQ3 - Summary (1) We summarize the domain-specific techniques used in previous research to apply LLMs to security tasks, including prompt engineering, fine-tuning, and external augmentation. (2) Prompt engineering is the most widely used domain technique, with almost all 127 papers employing this approach. Fine-tuning techniques were used in 25.2% of the papers, while task-specific external augmentation techniques were adopted in 22.05% of the papers. (3) We categorize and discuss the fine-tuning and external augmentation techniques mentioned in these papers, and analyze their relevance to specific security tasks.
@@ 8.1. Challenges
The generalization capability of LLMs pertains to their ability to consistently and accurately execute tasks across diverse tasks, datasets, or domains beyond their training environment. Despite undergoing extensive pre-training on large datasets to acquire broad knowledge, the absence of specialized expertise can present challenges when LLMs encounter tasks beyond their pre-training scope, especially in the cybersecurity domain. As discussed in Section  3, we explored the utilization of LLMs in 21 security tasks spanning five security domains. We observed substantial variations in the context and semantics of code or documents across different domains and task specifications. To ensure LLMs demonstrate robust generalization, meticulous fine-tuning, validation, and continuous feedback loops on datasets from various security tasks are imperative. Without these measures, there’s a risk of models overfitting to their training data, thus limiting their efficacy in diverse real-world scenarios.
@@ 9. Conclusion
LLMs are making waves in the cybersecurity field, with their ability to tackle complex tasks potentially reshaping many cybersecurity practices and tools. In this comprehensive literature review, we delved into the emerging uses of LLMs in cybersecurity. We firstly explored the diverse array of security tasks where LLMs have been deployed, highlighting their practical impacts (RQ1). Our analysis covered the different LLMs employed in security tasks, discussing their unique traits and applications (RQ2). Additionally, we examined domain-specific techniques for applying LLMs to security tasks (RQ3). Lastly, we scrutinized the data collection and preprocessing procedures, underlining the importance of well-curated datasets in effectively applying LLMs to address security challenges (RQ4). We outlined key challenges facing LLM4Security and provided a roadmap for future research, outlining promising avenues for exploration.